OpenStudy (anonymous):

I'm having trouble with order of operations. Solve each the following equation for x: 2x=-2[5-2(2x-12)]

OpenStudy (anonymous):

alright here we go...

OpenStudy (anonymous):

lmao it's probably something simply i'm missing

OpenStudy (anonymous):

i got half off on a test for missing order of operations...

OpenStudy (anonymous):

i distributed the -2 to the 2x and -12 first...

OpenStudy (anonymous):

i got 2x=-10-4x+24 first. which i guess is wrong. lol

OpenStudy (angela210793):

2x=-2[5-2(2x-12)] 2x=-2[5-4x+24] 2x=-2(-4x+29) 2x=8x-58 58=6x x=58/6

OpenStudy (anonymous):

ohhhhhhhhh. wow

OpenStudy (anonymous):

angela is way faster at typing than me lol but hers is correct

OpenStudy (anonymous):

haha. thank you guys... i seemed to have not been taught this integral part in my lower level classes. :(

OpenStudy (angela210793):

:)

OpenStudy (anonymous):

it won't let me select correct answer for some reason. can you guys hook each other up with one? thanks again!!!